OpenStudy (anonymous):

find derivative: f(x)= ln (3x+1)^4/(2x-1)^3

OpenStudy (anonymous):

is ln for the entire fraction or just for (3x+1)^4?

OpenStudy (anonymous):

\[\ln \frac{ (3x+1)^{4} }{(2x-1)^{3} } = \ln(3x+1)^{4}- \ln(2x-1)^{3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

differentiate each one individually in the form of \[\frac{ f'(x) }{ f(x) }\]

OpenStudy (anonymous):

and apply chain rule for each f'(x)

OpenStudy (anonymous):

can you do it now?

OpenStudy (anonymous):

im trying it now....thanks

OpenStudy (anonymous):

\[\ln (3x+1)^4/\ln (2x-1)^3= \ln (3x+1)^4-\ln (2x-1)^3= 4*3\ln(3x+1)^3-2*2\ln (2x-1)^2 \]